Hagenberg Sequent Symmetry Transputer

RISC

Linz

nCUBE 2

Salzburg

Convex Transputer Wien Intel Hypercube iPSC/860 Sequent Balance TU-Vienna

Wien

MTM-SUN Transputer Module

Multi- Cluster Transputer Module UNI-Vienna

[figure] [figure]

UNI-Salzburg GUP Intel Hypercube iPSC/860 MasPar MP- 1101A

PPS

Wien

64 kb/s

64 kb/s 64 kb/s